)A glocer mixes two kinds of nuts that costs $2.49 per pound and $3.89 per pound respectively, to make a 100 pounds of mixture that sosts $3.19 per pound. How much of each kind of nut is put into the mixture?

Answer by checkley75(3666)

100*319=249*X+389(100-X)
31900=249X+38900-389X
31900-38900=249X-389X
-7000=-140X
X=-7000/-140
X=50 POUNDS OF $2.49 NUTS.
100-50=50 POUNDS OF $3.89 NUTS.